# Depth-Width Tradeoffs in Algorithmic Reasoning of Graph Tasks with Transformers

### v1

@@ Abstract
Transformers have revolutionized the field of machine learning. In particular, they can be used to solve complex algorithmic problems, including graph-based tasks. In such algorithmic tasks a key question is what is the minimal size of a transformer that can implement a task. Recent work has begun to explore this problem for graph-based tasks, showing that for sub-linear embedding dimension (i.e., model width) logarithmic depth suffices. However, an open question, which we address here, is what happens if width is allowed to grow linearly. Here we analyze this setting, and provide the surprising result that with linear width, constant depth suffices for solving a host of graph-based problems. This suggests that a moderate increase in width can allow much shallower models, which are advantageous in terms of inference time. For other problems, we show that quadratic width is required. Our results demonstrate the complex and intriguing landscape of transformer implementations of graph-based algorithms. We support our theoretical results with empirical evaluations.
Machine Learning, ICML
@@ 1 Introduction
The transformer architecture (Vaswani, 2017), which was initially introduced for machine translation (Bahdanau, 2014), has emerged as the state-of-the-art neural network architecture across many fields, including computer vision (Dosovitskiy et al., 2021) and molecular analysis (Jumper et al., 2021). In order to explain how transformers emerged as the dominant neural architecture, it is important to understand the fundamental algorithmic primitives that transformers can solve in an efficient manner.
To that end, this paper addresses the following question.
Which architectural properties are necessary for transformers to solve graph algorithmic tasks?
In this work, we consider transformers that compute properties such as connectivity or the existence of a subgraph on some input graphs. We use this to develop a hierarchy of algorithmic tasks that proves sharp thresholds in the model width.
Graphs provide a compelling testbed for understanding transformer reasoning for serveral reasons. First, they serve as a natural “algorithmic playground,” encompassing a wide range of well-known problems that span computational classes. Many of these tasks have already been investigated as benchmarks for language models (Fatemi et al., 2023). Second, graph neural network (GNN) architectures (Gilmer et al., 2017; Kipf & Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017) are a well-established class of models for graph-structured data, but they have well-known limitations on their expressive power (Xu et al., 2018; Loukas, 2019). This raises an important question about whether transformers are a remedy for those structural limitations of GNNs.
Finally, inference-time reasoning in large language models (LLMs), such as chain-of-thought (Wei et al., 2022) and graph-of-thought (Besta et al., 2024) prompting can often be framed as a graph search over some solution space. By investigating transformers through the lens of graph algorithms, we aim to deepen our understanding of their algorithmic capabilities and the role of design choices, such as model width and depth.
Questions about transformer algorithmic capabilities have attracted considerable attention, and several results that characterize what can and cannot be calculated using transformers of a certain size have emerged (e.g. Merrill & Sabharwal, 2023b; Liu et al., 2023). However, we are still far from a complete understanding of the interplay between architecture and algorithmic capabilities. In this work we focus on this problem in the context of graph algorithms, with emphases on the role of model depth and width and on a particular node-adjacency tokenization scheme.
Recent works (Sanford et al., 2024b, c) developed trade-offs between the abilities of transformers of different depths to solve fundamental graph problems.
They identified parallelism as a key aspect that distinguishes transformers, proving surprising benefits of employing a transformer rather than a GNN for graph-based tasks. For example, while GNNs require depth $O(n)$ to determine whether an input graph with $n$ vertices is connected, transformers can solve this task using only $O(\log n)$ layers of self-attention.
However, these results rely on restrictive assumptions about how graphs are represented as inputs to neural models. First, they tokenize an input graph as a sequence of discrete edge embeddings, rather than more standard node-based embeddings. Second, they focus on the scaling regime where the transformer’s embedding dimension (which we refer to interchangeably as its width) is restricted to be much smaller than the size of the input graph.
While a small width is certainly desirable, practical transformers have large widths with strict limitations on model depth.
For example, when learning over graphs, the width of the networks is often much larger than the number of vertices in the graph. Table 3 shows a list of commonly used graph datasets, including many molecular datasets, where the average graph size in the data is less than $40$. This is while commonly used network width is above $64$.
In this paper, we seek to understand which graph algorithms can be implemented in realistic transformer scaling regimes, such as when the width grows linearly in the number of tokens and the depth is constant, independent of the graph size. Concretely, we consider transformers that take as input graphs with adjacency embeddings for each node and place upper and lower bounds on the minimum width needed to solve several tasks. We focus on the constant-depth regime, where we do not allow the number of attention layers to scale as a function of the input size.
In Section 4, we introduce tasks—such as graph connectivity and fixed-length cycle detection—for which linear width is necessary and sufficient for dense graph inputs. Bounded-degree assumptions reduce the width threshold for the cycle detection task. Section 5 shows that more complex tasks—such as subgraph counting and Eulerian cycle verification—require super-linear width. We visualize the hierarchy over transformer widths induced by our collection of positive and negative results in Figure 1, which ranges from local node-level tasks that can be solved with constant width (such as computing the degree of each node), to arbitrary functions that necessitate a quadratic width scaling.
@@
Figure 1: The width complexity hierarchy of graph tasks for transformers with node-adjacency tokenizations and constant depth. Each row visualizes the width regimes where the task is solvable (blue) or hard (red).
We validate our results empirically in Section 6. Specifically, we show that transformers with large embedding dimension and low depth achieve similar accuracy to deeper models, while having faster training and inference time, due to GPU parallelization.
@@ 2 Related Works
Transformer with arbitrary depth (Wei et al., 2021; Yun et al., 2020) or arbitrarily many chain-of-thought tokens (Malach, 2023) are known to be universal approximators. While these universality results are fundamental, they leave open questions about the ability of transformers to solve tasks efficiently, particularly in terms of parameter count and depth. To address this, various theoretical techniques have been employed to develop a more precise understanding of the scaling trade-offs involved.
@@
For instance, Merrill & Sabharwal (2023a) show that constant-depth transformers with polynomial width can be simulated by $\mathsf{TC^{0}}$ Boolean circuits, implying that they cannot solve problems like graph connectivity. These results provide essential context for our paper; while we prove that transformers with a constant number of attention layers and bounded width can solve such tasks (e.g. in Theorem 4.1), this does not contradict their bounds, and together, they imply that a sufficiently large depth or width is necessary in the multi-layer perceptrons of our constructions. Similarly, Hao et al. (2022) identify formal language classes that can and cannot be recognized by hard-attention transformers. On the other hand, deeper transformers or those with sufficiently long chain-of-thought reasoning can simulate finite-state automata (Liu et al., 2023; Li et al., 2024).
@@
A different perspective frames transformers in terms of communication complexity. Sanford et al. (2024c, b) draw an analogy between transformers and the Massively Parallel Computation (MPC) framework (Karloff et al., 2010), similar to prior work linking GNNs to the Congest model in distributed computing (Loukas, 2019). Sanford et al. (2024a) extend this analogy to define a transformer complexity hierarchy for graph tasks. While they focus on an edge-list tokenization and depth-based hierarchy, our work considers a width-based hierarchy with a node-adjacency tokenization, leading to different complexity insights.
Graph transformers.
Recent advancements in graph transformer architectures have significantly expanded the capabilities of Graph Neural Networks (GNNs) (Gilmer et al., 2017) by leveraging the power of attention mechanisms (Veličković et al., 2018; Brody et al., 2022) to facilitate dynamic node aggregation based on neighboring structures without relying on positional encodings. Building on this foundation, models like GraphBert (Zhang et al., 2020) incorporated graph-structured node embeddings to enhance representation learning. The evolution continued with the development of comprehensive graph transformer models that extend beyond traditional message-passing paradigms. For instance  Kreuzer et al. (2021) proposed graph transformers that generalize sequence-based transformers Subsequent models, including Spectral Attention Networks Kreuzer et al. (2021), introduced novel mechanisms for integrating spectral information and enhancing scalability for larger graphs. Additionally, GraphGPS (Rampášek et al., 2023) exemplifies the hybrid approach by combining attention mechanisms with message-passing techniques, thereby harnessing the strengths of both methods.
There is an abundance of possible embedding for graph transformers (Kreuzer et al., 2021; Dwivedi & Bresson, 2021; Kim et al., 2022; Rampášek et al., 2023; Zhang et al., 2024). Our work does not focus on comparing between such embeddings, but rather to study theoretically the effect of the embedding dimension. The embedding type itself is mainly a technical detail for proving our theoretical guarantees.
@@ 3.1 Transformers
We consider the following setting of transformers: The input is a sequence of $N$ tokens $\mathbf{x}_{1},\dots,\mathbf{x}_{N}$ where $\mathbf{x}_{i}\in\mathbb{R}^{d_{\text{in}}}$. We denote by $X^{(0)}\in\mathbb{R}^{d_{\text{in}}\times N}$ the matrix where each row $i$ is equal to $\mathbf{x}_{i}$. Each layer of the transformer applies a self-attention mechanism on the inputs, and then an MLP. We denote by the input to layer $\ell$ as $X^{(\ell-1)}$. The self-attention at layer $\ell$ with $H$ heads is parameterized by matrices $K^{(\ell)}_{h},Q_{h}^{(\ell)}\in\mathbb{R}^{m_{\ell-1}\times m_{\ell-1}}$, $V^{(\ell)}_{h}\in\mathbb{R}^{m_{\ell-1}\times m_{\ell-1}}$. It is defined as:
$Z^{(\ell)}=\sum_{h=1}^{H}V^{(\ell)}_{h}X^{(\ell-1)}\text{sm}(X^{(\ell-1)^{\top}}K_{h}^{(\ell)^{\top}}Q^{(\ell)}_{h}X^{(\ell-1)})~{},$
where sm is row-wise softmax. The output is of dimension $Z^{(\ell)}\in\mathbb{R}^{m_{\ell-1}\times N}$. After that we use a residual connection, so the output of the self-attention layer is $\tilde{X}^{(\ell)}=Z^{(\ell)}+X^{(\ell-1)}$ Finally, we apply an MLP $\mathcal{N}^{(\ell)}:\mathbb{R}^{m_{\ell-1}}\rightarrow\mathbb{R}^{m_{\ell}}$ with ReLU activations on each token separately (i.e. each column of $\tilde{X}^{(\ell)}$). The output of the MLP is denoted as $X^{(\ell)}$, which is the input to the next layer. We denote by $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ the ReLU activation.
We also often use positional encodings which are commonly used in standard transformer architectures. It is also common to add normalization layers, however it will not be necessary for our results, and these layers can be trivially added in all of our results without changing the output. The bit-precision of all our transformers will be $O(\log(n))$, where $n$ is the number of input tokens. This is a common assumption in many previous works (Sanford et al., 2024c, b; Merrill & Sabharwal, 2023b, 2024), and is relatively mild. We will denote $m=\max(m_{0},\dots,m_{L})$ the embedding dimension, and $d_{\text{in}}$ the input dimension.
@@
Unlike GNNs, transformers are sequential architectures whose network topology does not encode the structure of an input graph. It is therefore a significant modeling decision to decide how a graph is best tokenized into a sequence of embeddings that can be proved to a transformer.
In this work, we focus primarily on the node-adjacency tokenization scheme, where each token corresponds to a node and encodes its incident edges.
Given a graph $G$ with $n$ nodes, let $A\in\mathbb{R}^{n\times n}$ be its adjacency matrix. The $i$th token input $\mathbf{x}_{i}\in\mathbb{R}^{n}$ to the transformer is defined as the $i$th row of $A$. Thus, $X\in\mathbb{R}^{n\times n}$.
There are numerous alternatives, but we restrict our focus to two others for experimental comparisons in Section 6. The edge-list tokenization of Sanford et al. (2024a) converts the graph into a sequence of discrete edge tokens. For a unique identifier assigned to each $v\in V$, this scheme encodes each edge $(v_{i},v_{j})\in E$ as a token $\mathbf{x}=(i,j)\in\mathbb{R}^{2}$.
A further alternative is the Laplacian eigenvector tokenization, which is introduced and discussed in Appendix A.
3.2 Graphs and Parallel Computation
Sanford et al. (2024b, a) introduced a representational equivalence between transformers with $L$ layers and $O(L)$ rounds of massively parallel computation (MPC). MPC (Karloff et al., 2010) is a theoretical formalization of the MapReduce (Dean & Ghemawat, 2008) distributed computing protocol. Under MPC, a very large number of machines with limited memory alternate between local computation and global communication rounds. The equivalence implies that transformers can act as MPC protocols by tokens as machines, local computation as MLP layers, and communication as self-attention. Critically, GNNs are not equivalent to MPC in the same sense, and are rather equivalent to the Local or Congest models, depending the sizes of each message between nodes is bounded (see Loukas, 2019). The main advantage of transformers over GNNs is the use of parallelism.
Graph algorithms are well-understood in the MPC framework. For instance, a wide range of graph tasks can be organized into depth-based equivalence classes that depend on the tasks’ computational complexity (Nanongkai & Scquizzato, 2022). While few unconditional negative results are known in the MPC framework, the one-cycle vs two-cycle task—which asks whether an input graph is either (1) a single cycle graph with $n$ nodes, or (2) a pair of cycle graphs, with $\frac{n}{2}$ nodes each—is widely expected to require either $\Omega(\log n)$ rounds or local memory of $\Omega(n)$. By combining these algorithms and conditional lower bounds with the transformer equivalence, we can obtain a numerous size complexity thresholds for transformers.
However, the results in Sanford et al. (2024b, a) have two limiting assumptions. The first is that the embedding dimension of the transformer
is strictly sublinear, namely $m=n^{1-\epsilon}$ for some constant $\epsilon\in(0,1)$ independent of $n$. The second limitation is that the graph is represented as an edge list, which requires $\Omega(n^{2})$ tokens for dense graphs. In this work we go beyond the strict sublinear regime, and focus on the node-adjacency tokenization strategy.
Remark 3.1.
When calculating the complexity of solving certain graph tasks, we usually only count the number of transformer layers, while the MLP can be arbitrarily large. This corresponds to the MPC model where usually only the distributed computing time is considered, while it is assumed that there is an unlimited local compute time for each machine. Intuitively, the number of self-attention layers is equivalent to the number of distributed computation rounds, while the MLP represent the local computation of each machine.
4 Beyond Sub-linear Embedding Dimension
In this section we show that having a transformer with linear embedding dimension can easily improve some known lower bounds based on distributed computing. We also show new lower bounds that are stricter for transformers with a linear embedding dimension, and a matching upper bound. All the proofs for the theorems in this section are in Appendix B.
4.1 Improving the sub-linear lower bounds

### v2

# Depth-Width Tradeoffs for Transformers on Graph Tasks

@@ Abstract
Transformers have revolutionized the field of machine learning. In particular, they can be used to solve complex algorithmic problems, including graph-based tasks. In such algorithmic tasks a key question is what is the minimal size of a transformer that can implement the task. Recent work has begun to explore this problem for graph-based tasks, showing that for sub-linear embedding dimension (i.e., model width) logarithmic depth suffices. However, an open question, which we address here, is what happens if width is allowed to grow linearly, while depth is kept fixed. Here we analyze this setting, and provide the surprising result that with linear width, constant depth suffices for solving a host of graph-based problems. This suggests that a moderate increase in width can allow much shallower models, which are advantageous in terms of inference and train time. For other problems, we show that quadratic width is required. Our results demonstrate the complex and intriguing landscape of transformer implementations of graph-based algorithms. We empirically investigate these trade-offs between the relative powers of depth and width and find tasks where wider models have the same accuracy as deep models, while having much faster train and inference time due to parallelizable hardware.
@@ 1 Introduction
The transformer architecture (Vaswani, 2017), has emerged as the state-of-the-art neural network architecture across many fields, including language (Brown et al., 2020), computer vision (Dosovitskiy et al., 2021) and molecular analysis (Jumper et al., 2021). This success prompts a basic question: which algorithms can be implemented using transformers, and how do specific architecture choices affect this capability. Understanding these questions could enable practitioners to design more computationally efficient models without compromising expressivity.
Two significant architectural elements of transformers are the depth (the number of layers) and width (the dimension of their latent representation). Increasing either property increases the expressivity of the underlying architecture, but the qualitative differences between “shallow and wide” and “deep and narrow” architectures are poorly understood. Given the widespread use of specialized parallel hardware for training and serving modern language models, increasing the width has a limited effect on model latency, provided sufficient memory. A wide range of recent theoretical work (see e.g., Merrill and Sabharwal, 2023b; Liu et al., 2023) explores these fundamental expressivity questions for various benchmark tasks and transformer scaling regimes. However, a complete understanding of the interplay between architecture and algorithmic capabilities—and in particular the fine-grained interplay between depth and width—remains elusive. In this work, we aim to crystallize the emergent algorithmic capabilities of transformers as a function of model depth and width, and demonstrate occasions when, contrary to conventional wisdom, the benefits of width eclipse those of depth. Specifically, we focus on the following question:
What are the algorithmic capabilities obtained by increasing the width of a transformer?
We ground this question in the context of graph algorithmic tasks, which provide a compelling testbed for understanding transformer reasoning. Graph algorithms serve as a natural “algorithmic playground,” encompassing a wide range of well-known problems that span computational classes. Many of these tasks have already been investigated as benchmarks for language models (Fatemi et al., 2023). Furthermore, previous theoretical works have routinely employed these algorithmic tasks to unveil the capabilities and limitations of graph neural networks (GNNs, Gilmer et al., 2017). For example, the theoretical analysis of Loukas (2019) revealed lower bounds on the depth and width necessary to determine connectivity. Sanford et al. (2024a) employed similar analysis to elucidate such trade-offs for transformers. However, their results pertain primarily to the powers of depth and a specific edge-list representation. Other scaling regimes—such as transformers with fixed depth and variable width—have yet to be explored in the context of this toolkit.
For such graph reasoning tasks, we are principally concerned with the optimal transformer size scaling as a function of the graph size. Specifically, given a graph with $n$ nodes and an encoding of an algorithmic task, such as determining connectivity or counting the number of triangles, we investigate the width necessary and sufficient to solve the task as a function of $n$. This fixed-depth, variable-width setting is arguably more pertinent given the aforementioned benefits of width in computational efficiency. Indeed, practical applications of transformers to graphs have width that is often much larger than the number of vertices in the graph. [^1] See Table 3 for a list of commonly used graph datasets (including molecular datasets) where the average graph size in the data is less than $40$.
Our theoretical contribution is a novel representational hierarchy that characterizes this depth-width dependence in transformers. Surprisingly, our results show that multiple problems of interest can be solved with fixed depth and linear width with respect to $n$. We provide empirical support for the advantage of these shallow and wide models.
@@
Figure 1: The width complexity hierarchy of graph tasks for transformers with constant depth. Each row visualizes the width regimes where the task is solvable (blue) or hard (red).
Our results elucidate a hierarchy concerning the minimum width needed for constant-depth transformers to represent solutions to various graph algorithmic tasks. Section˜5 establishes the relevance of the linear scaling regime, where the width of the transformer increases linearly in the number of nodes $n$. Specifically, it considers tasks, such as graph connectivity and detection of fixed length cycles, for which linear width is necessary and sufficient for dense graph input. Section˜5 demonstrates that more complex tasks, including subgraph counting and Eulerian cycle verification, require super-linear and near-quadratic width respectively. The resulting hierarchy over transformer widths induced by our collection of positive and negative results is visualized in Figure˜1, which ranges from local node-level tasks that can be solved with constant width (such as computing the degree of each node), to arbitrary functions that require quadratic width scaling. Some of the lower bounds are conditional on the well known one vs. two cycles conjecture (see section˜B.1 for a formal statement).
The empirical results of Section˜6 demonstrate the relevance of our results in practice. We consider various graph problems, and experiment with varying the depth and width of the model under a constraint on the overall parameter count. Our results demonstrate that shallower and wider models are significantly faster than narrow and deep models.
@@ 2 Related Works
Transformer with arbitrary depth (Wei et al., 2021; Yun et al., 2020) or arbitrarily many chain-of-thought tokens (Malach, 2023) are known to be universal approximators. Informally, these results suggest that any algorithm can be simulated with a sufficiently large transformer. They, however, leave open more fine-grained questions about more compact transformer implementations of a given algorithm, particularly in terms of parameter count and depth.
@@
Various theoretical techniques have been employed to develop a more precise understanding of the relevant scaling trade-offs. For instance, Merrill and Sabharwal (2023a) show that constant-depth transformers with polynomial width can be simulated by $\mathsf{TC^{0}}$ Boolean circuits, implying that they cannot solve problems like graph connectivity. These results provide essential context for our paper, where we introduce scaling regimes where transformers can and cannot learn such tasks.Similarly, Hao et al. (2022) identify formal language classes that can and cannot be recognized by hard-attention transformers. Another way to augment transformer expressivity is to use chain-of-thought reasoning (Wei et al., 2022). Indeed, it has been shown that sufficiently long chains reasoning can simulate finite-state automata (Liu et al., 2023; Li et al., 2024).
@@
A different perspective frames transformers in terms of communication complexity. Sanford et al. (2024c, b) draw an analogy between transformers and the Massively Parallel Computation (MPC) framework (Karloff et al., 2010), similar to prior work linking GNNs to the Congest model in distributed computing (Loukas, 2019). Sanford et al. (2024a) extend this analogy to define a transformer complexity hierarchy for graph tasks.
Graph transformers and GNNs
Research in transformers and graph neural networks (GNNs) reveals a rich overlap, primarily because both are fundamentally message-passing models. Substantial research effort has been devoted to the study of hybrid architectures to combine GNNs’ intrinsic encoding of input graphs and transformers’ empirical successes. A variety of approaches to model hybridization have been explored, including the incorporation of graph structure directly into attention layers (Veličković et al., 2018; Brody et al., 2022), as subgraph features in positional encodings (Zhang et al., 2020), and as spectral features of the graph Laplacian matrix (Kreuzer et al., 2021). The aim of this work is to understand transformer architectural trade-offs rather than to introduce a novel architecture. However, the graph encoding schemes employed in this paper draw upon key insights from this literature. We primarily utilize a simple adjacency encoding in order to isolate the focus of our study to the impacts of changing the embedding dimension, but we also investigate some theoretical and empirical trade-offs for spectral Laplacian-based encodings (Appendix˜A).
@@ 3.1 Transformers
We consider the following setting of transformers: the input is a sequence of $N$ tokens $\mathbf{x}_{1},\dots,\mathbf{x}_{N}$ where $\mathbf{x}_{i}\in\mathbb{R}^{d_{\text{in}}}$. We denote by $X^{(0)}\in\mathbb{R}^{d_{\text{in}}\times N}$ the matrix where the $i$-th column is equal to $\mathbf{x}_{i}$. Each layer of the transformer applies a self-attention mechanism on the inputs, and then an MLP. We denote the input to layer $\ell$ by $X^{(\ell-1)}$. The self-attention at layer $\ell$ with $H$ heads is parameterized by matrices $K^{(\ell)}_{h},Q_{h}^{(\ell)}\in\mathbb{R}^{m\times m}$, $V^{(\ell)}_{h}\in\mathbb{R}^{m\times m}$. [^2] Except for the first layer where $K^{(\ell)}_{h},Q_{h}^{(\ell)}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{in}}}$, $V^{(\ell)}_{h}\in\mathbb{R}^{m\times d_{\text{in}}}$ It is defined as:
$Z^{(\ell)}=\sum_{h=1}^{H}V^{(\ell)}_{h}X^{(\ell-1)}\textsc{sm}(X^{(\ell-1)^{\top}}K_{h}^{(\ell)^{\top}}Q^{(\ell)}_{h}X^{(\ell-1)})~,$
where sm is row-wise softmax, and $m$ is the hidden embedding dimension. The output is of dimension $Z^{(\ell)}\in\mathbb{R}^{m\times N}$. This is followed by a residual connection, so that the output of the self-attention layer is $\tilde{X}^{(\ell)}=Z^{(\ell)}+X^{(\ell-1)}$.
Finally, we apply an MLP $\mathcal{N}^{(\ell)}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ with ReLU activations on each token separately (i.e. each column of $\tilde{X}^{(\ell)}$). The output of the MLP is $X^{(\ell)}$, and is the input to the next layer. We let $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ denote the ReLU activation. Critically, we assume that the MLP layer can compute arbitrary functions of each individual token embedding $X^{(\ell)}_{i}$. While involved, this assumption is common in the theoretical literature (see e.g. Sanford et al., 2024b) and enables a theoretical understanding of the modeling limitations imposed by the attention layer. However, we note that in all our constructions, the size of the MLP is at most polynomial in the number of nodes in the input graph. Our MLP layers further incorporate arbitrary positional encodings of the index $i$ of each token embedding $\tilde{X}_{i}^{(\ell)}$, i.e. $\mathcal{N}^{\ell}(\tilde{X}^{(\ell)})=(g(\tilde{X}_{1}^{(\ell)},1),\dots,g(\tilde{X}_{N}^{(\ell)},N)),$ for some $g$. This formalism allows standard transformer positional encodings to be implemented in our theoretical model. Normalization layers—which are not explicitly accounted for in our transformer definition—can be similarly incorporated by taking advantage of the arbitrary MLP assumption without changing the results.
The bit-precision of all our transformers will be $O(\log n)$, where $n$ is the number of input tokens. This is a common assumption in many previous works Sanford et al. (2024c, b); Merrill and Sabharwal (2023b, 2024), and is relatively mild. It is also a necessary requirement for representing a number of size $n$, e.g. when having positional embeddings for $n$ tokens. We will denote $m=\max(m_{0},\dots,m_{L})$ the embedding dimension, and $d_{\text{in}}$ the input dimension.
@@
Because the network topology of transformers (unlike GNNs) does not encode the structure of an input graph, graph structure must be encoded explicitly into the input of the transformer. Choosing the format of this tokenization scheme for graph inputs is a significant modeling decision. The most fundamental aspect of that choice involves deciding between edge-wise and node-wise schemes. Sanford et al. (2024a) employ an edge-list tokenization that converts the graph into a sequence of discrete edge tokens. However, this scheme suffers computationally for large graphs; the quadratic computational bottleneck of self-attention can result in an $\Omega(n^{4})$ runtime for dense graphs with $n$ nodes.
In contrast, we consider node-wise encodings, where each node corresponds to exactly one input embedding. Our primary theoretical results use what is arguably the simplest such representation: the node-adjacency tokenization scheme. In this representation, each input embedding directly encodes a node’s edges as an adjacency vector. That is, for graph $G$ with $n$ nodes and adjacency matrix $A\in\{0,1\}^{n\times n}$, the $i$th token input $\mathbf{x}_{i}\in\mathbb{R}^{n}$ to the transformer is defined as the $i$th row of $A$. An alternative node-wise scheme is the Laplacian eigenvector tokenization (Kreuzer et al., 2021), which captures global graph structure node-wise in a lossy manner with the most significant eigenvectors of the Laplacian matrix. We empirically contrast the node-adjacency scheme with both alternatives in Section˜6, and we further explore the representational properties of the Laplacian tokenization in Appendix˜A.
The adjacency node embedding is not permutation invariant, similar to the edge-list embedding used in Sanford et al. (2024a). Namely, if the tokens contain positional embeddings, then changing the order of the tokens may also change the output. These embeddings are used for technical convenience, especially since spectral embeddings that are permutation invariant are difficult to analyze for combinatorial problems (such as connectivity, subgraph counting, etc.). In section˜6 we provide experiments that justify the use of such embedding on real-world datasets.
4 The Expressive Power of a Linear Embedding Dimension
In this section we focus on two graph reasoning tasks: the $1$ vs. $2$ cycle problem and the cycle detection problem. We show that a fixed-depth transformer can solve these tasks with only linear width. We also show that these results are tight. All the proofs for the theorems in this section are in Appendix˜B. Finally, in Sec. 4.3 we show that for cycle detection in bounded degree graphs, sublinear width suffices. Note that the node degree task as appears in fig.˜1 (namely, calculating the degree of each node) can be easily solved with an $O(\log(n))$ embedding size, as shown e.g., in Sanford et al. (2024a)
4.1 A linear width solution for $1$ vs. $2$ cycle detection